OpenStudy (anonymous):

the point (8,-15) is on the terminal side of an angle pie.what is sin pi?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(8,-15) = (x, y) \[r^2 = x^2+y^2 \rightarrow r^2 = (8)^2+(-15)^2\] \[\rightarrow r^2 = 64 + 225 = 289 \rightarrow r = 17\] ( where r is the hypotenuse) Now, \[\sin \pi = \frac {y}{r} \rightarrow \sin \pi = \frac {-15}{17}\] @zebulin

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

help wiith another?

OpenStudy (anonymous):

@zebulin medal please